OpenStudy (anonymous):

Use the value of the trigonometric function to evaluate the indicated function. Sin t= 1/2 Csc (-t)=?

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ sin(-\theta)\implies -sin(\theta)}} \\ \quad \\ \quad \\ sin(t)=\cfrac{1}{2}\qquad \qquad csc(-t)\implies \cfrac{1}{sin(-t)}\) what do you think?

OpenStudy (anonymous):

-0.5???

OpenStudy (jdoe0001):

well close \(\bf {\color{blue}{ sin(-\theta)\implies -sin(\theta)}} \\ \quad \\ \quad \\ sin(t)=\cfrac{1}{2}\qquad \qquad csc(-t)\implies \cfrac{1}{sin(-t)}\implies \cfrac{1}{-sin(t)} \implies \cfrac{1}{-\frac{1}{2}} \\ \quad \\ recall \implies \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\qquad thus \\ \quad \\ \cfrac{1}{-\frac{1}{2}}\implies \cfrac{1}{1}\cdot -\cfrac{2}{1}\)